### v1

# From Prompts to Protection: Large Language Model-Enabled In-Context Learning for Smart Public Safety UAV

Abstract
A public safety Unmanned Aerial Vehicle (UAV) enhances situational awareness in emergency response. Its agility and ability to optimize mobility and establish Line-of-Sight (LoS) communication make it increasingly vital for managing emergencies such as disaster response, search and rescue, and wildfire monitoring. While Deep Reinforcement Learning (DRL) has been applied to optimize UAV navigation and control, its high training complexity, low sample efficiency, and simulation-to-reality gap limit its practicality in public safety. Recent advances in Large Language Models (LLMs) offer a compelling alternative. With strong reasoning and generalization capabilities, LLMs can adapt to new tasks through In-Context Learning (ICL), which enables task adaptation via natural language prompts and example-based guidance, without retraining. Deploying LLMs at the network edge, rather than in the cloud, further reduces latency and preserves data privacy, thereby making them suitable for real-time, mission-critical public safety UAVs. This paper proposes the integration of LLM-enabled ICL with public safety UAV to address the key functions, such as path planning and velocity control, in the context of emergency response. We present a case study on data collection scheduling where the LLM-enabled ICL framework can significantly reduce packet loss compared to conventional approaches, while also mitigating potential jailbreaking vulnerabilities. Finally, we discuss LLM optimizers and specify future research directions. The ICL framework enables adaptive, context-aware decision-making for public safety UAV, thus offering a lightweight and efficient solution for enhancing UAV autonomy and responsiveness in emergencies.
Index Terms:
Unmanned Aerial Vehicle, Large Language Models, In-Context Learning, Network Edge
I Introduction
Unmanned Aerial Vehicles (UAVs) play a crucial role in various sectors, including energy, environmental monitoring, parcel delivery, and public safety. In particular, they are transformative in public safety applications, including post-disaster assessment, search and rescue missions, wildfire monitoring, border surveillance, and crime scene reconstruction, due to their agility, optimized mobility, and ability to establish reliable Line-of-Sight (LoS) communications. A public safety UAV enhances situational awareness for police, firefighters, and other first responders, thereby improving incident management and overall operational efficiency[1],[2].
UAVs can leverage Machine Learning (ML) techniques for optimizing navigation and control, particularly in emergencies. For example, they employ Deep Reinforcement Learning (DRL) to optimize flight trajectories based on victims’ needs or to identify casualties and collect environmental data efficiently[3]. However, DRL solutions involve complex training processes and face challenges such as the simulation-to-reality gap and low sample efficiency, which are ill-suited to the urgent demands of emergency scenarios. Therefore, lightweight and adaptive solutions are required to quickly adapt to the emergency and enhance the success rate of rescue operations.
Recent advancements in artificial intelligence, particularly in Large Language Models (LLMs), are driving significant transformations across UAV communications[4]. With strong reasoning and generalization capabilities, LLMs enable advanced understanding, flexible adaptation, and real-time responsiveness in diverse UAV scenarios. Integrating LLMs into UAV communications offers a promising path toward greater autonomy, equipping UAVs with enhanced decision-making abilities and enabling more effective responses during emergencies.
The conventional approach to deploying LLMs relies on cloud-based infrastructures, which often introduces communication delays and raises privacy concerns, particularly problematic for time-sensitive and mission-critical UAV operations[5]. A recent trend is to shift LLM deployment to the network edge through edge server integration. This approach allows LLMs to directly serve the computational and operational needs of edge devices such as UAVs. Notably, LLMs can be taught new tasks using In-Context Learning (ICL) by the UAV, enabling rapid adaptation without retraining. For instance, ICL can facilitate key UAV functions, including path planning and velocity control, data collection scheduling, and power management, enhancing real-time decision-making in emergencies.
Figure 1: Illustration of LLM-enabled ICL for a public safety UAV. The public safety UAV collects sensory data and submits it to the edge LLM, where a task description, which is a basis for ICL, is formulated with the following components: task goal, input data, rules, output, and feedback. The system prompt defines the rule of LLM, e.g., data collector.
This paper proposes the use of LLM-enabled ICL to facilitate the key functions in public safety UAV. We systematically categorize these functions and present detailed task descriptions for each, demonstrating how ICL can enable the public safety UAV to adapt rapidly to emergency scenarios, improve decision-making, and enhance operational efficiency without requiring model retraining. As depicted in Fig. 1, ICL leverages structured natural language prompts, i.e., task descriptions, to encapsulate task objectives, while curated demonstration sets convey relevant system states and environmental conditions such as sensory data. This approach enables LLMs to infer an effective output for the task, without requiring explicit model retraining. By using contextually rich examples, ICL supports adaptive, on-the-fly decision-making across a broad range of UAV-Assisted Sensor Network (UASNETs) tasks, making it particularly valuable for time-sensitive applications such as emergency response.
The contributions of the paper are listed as follows:
  • A classification of key functions in public safety UAVs, such as path planning and velocity control. These functions can be effectively addressed using LLM-enabled ICL, providing a structured foundation for its application in emergencies. This classification serves as a foundational reference to guide future research and development in integrating LLM-enabled ICL into public safety UAVs, especially where DRL optimizers fall short.
  • A case study on data collection schedule, showcasing how an LLM-enabled ICL framework significantly reduces packet loss compared to conventional strategies, while also addressing potential jailbreaking vulnerabilities.
  • A discussion on LLM optimizers along with future research directions. One such direction involves the use of Diffusion Models (DMs) to generate diverse, high-quality synthetic examples that can be used for enriching task descriptions.
The remainder of this paper is organized as follows: Section II discusses the deployment of LLMs. Section III provides background on LLM-enabled ICL and its effectiveness. Section IV explores the application of LLM-enabled ICL for public safety UAV and its various functions. Section V presents a case study on data collection schedule, along with investigating jailbreaking attacks. Section VI discusses LLM optimizers and their merits for public safety UAV. Section VII outlines future directions, including the use of DMs to enrich task descriptions. Finally, Section VIII concludes the paper.
Table I: Overview of In-Context Learning for LLMs in Public Safety UAV Applications.
Aspect Description
Definition ICL enables LLMs to learn from contextual information embedded in prompts containing task-specific instructions and demonstrations written in natural language.
Capabilities Allows LLMs to generalize across tasks by leveraging pre-trained knowledge and adapting to new tasks through example-based prompts. Considered a core feature of modern LLMs.
Key Considerations The design of demonstrations, including example selection, formatting, and ordering, strongly influences performance and should be carefully crafted.
Application in Public Safety ICL offers a lightweight method to facilitate key functions such as path planning by reusing prior solutions as input prompts, supporting fast adaptation in emergencies.
II Deployment of LLMs
This section highlights the importance of LLM deployment and discusses limitations of cloud-based deployment and the feasibility of edge-based LLM deployment. LLMs impose substantial demands on computational and storage resources. For example, GPT-4 contains approximately 1.76 trillion parameters and has a model size of 45 GB, placing significant strain on network storage infrastructure. Fine-tuning even a comparatively smaller LLM, such as a 7-billion-parameter model like GPT4-LLM, can require nearly three hours on a high-end setup with 8×80GB A100 GPUs—making the process highly time-intensive. Moreover, LLM inference contributes to overall network latency, influenced by factors such as hardware configuration, batch size, degree of parallelism, and optimization techniques like model pruning. Therefore, efficient and strategic deployment of LLMs is essential to meet the performance and scalability demands of UASNETs[6].
Thanks to their strong generalization and reasoning capabilities, LLMs enable UAVs to interpret complex, dynamic environments and develop coherent, sequential control strategies for navigation, surveillance, and decision-making tasks. These capabilities make LLMs well-suited for real-time operations in mission-critical scenarios such as disaster response, infrastructure inspection, and surveillance. However, most existing LLM deployments rely heavily on cloud-based infrastructures, which impose critical limitations in terms of latency, network bandwidth, and data privacy, making them suboptimal or even unfeasible for delay-sensitive or communication-limited UAV applications. To address this problem, efficient model compression techniques such as quantization, pruning, and knowledge distillation are used to minimize model size and computational load without compromising performance. In parallel, fast decoding strategies such as speculative decoding, early termination, parallel decoding, and sparse attention reduce inference latency and memory requirements during autoregressive token generation.
Figure 2: Task Descriptions for key functions in public safety UAV comprising task goal, input data, rules and constraints, output, and feedback loop.
Once LLMs are efficiently deployed at the edge, they can be leveraged for context-aware inference through prompt engineering. Techniques such as CoT prompting, prompt-based planning, self-refinement prompting, and ICL enable LLMs to perform complex tasks by interpreting instructions and demonstrations embedded in prompts without the need for parameter updates. ICL is particularly effective in generalizing tasks because it uses previously trained knowledge and adapts to examples, with performance depending on the design of the prompts, including the selection, arrangement, and formatting of the demonstrations. In dynamic, resource-constrained scenarios such as UASNETs, ICL offers tangible benefits as previous solutions can serve as reusable inputs for new challenges, increasing the flexibility and responsiveness of UAVs. [7].
III LLM-enabled In-Context Learning
This section provides background on LLM-enabled ICL and discusses its effectiveness. ICL refers to the ability of a pretrained language model to perform a task by conditioning on a sequence of input-output demonstration examples, without requiring any updates to the model parameters. Given an input query x and a set of candidate answers $\mathcal{Y}=\{y_{1},\ldots,y_{m}\}$, the model selects the output with the highest likelihood based on a context $C$, which includes an optional task instruction I and k demonstration examples. This context can be expressed as either:
$\mathcal{C}=\{I,s(x_{1},y_{1}),\ldots,s(x_{k},y_{k})\}$
where $s$ denotes a natural language representation of the examples.
ICL differs from related learning paradigms in the following ways:
  1. 1.
    Prompt Learning: While prompt learning uses fixed textual templates or soft prompt embeddings to elicit model behavior, ICL can be regarded as a subclass of prompt learning where the demonstrations are included directly in the task description.
  2. 2.
    Few-Shot Learning: Traditional few-shot learning involves adapting the model’s parameters using a small number of supervised examples. In contrast, ICL requires no parameter updates and leverages the pretrained model directly through inference over the provided context[8]
Figure 3: Illustration of a public safety UAV, where the UAV follows the trajectory and establishes LoS communication with the ground sensor and uses the LLM as an optimizer.
The likelihood of a candidate answer $y_{j}$ is computed using a scoring function $f$ is given by:
$P(y_{j}\mid x)\triangleq f_{M}(y_{j},\mathcal{C},x),$
and the final prediction $\hat{y}$ is given by:
$\hat{y}=\arg\max_{y_{j}\in\mathcal{Y}}P(y_{j}\mid x).$
Min et al.[9] reveal that the effectiveness of ICL relies more on the structural presentation of input and label spaces than on the correctness of individual demonstration labels. Surprisingly, replacing ground truth labels with random ones in demonstrations has only a marginal effect on performance. Moreover, models can retain up to 95% of performance gains using either inputs or label sets alone—provided the formatting is appropriate. These insights suggest that formatting and representational cues play a more critical role than previously assumed, with implications for both model design and future research in prompt engineering and meta-training for ICL. Table I provides an overview of LLM-enabled ICL for public safety UAV. The key point is that the way demonstrations are designed, such as the choice of examples, their formatting, and sequence, has a significant impact on performance and must be meticulously crafted.
IV LLM-enabled In-Context Learning for Optimized Public Safety UAV
This section highlights the potential of LLM-enabled ICL in optimizing key functions in public safety to dynamically adapt based on real-time context and prior experiences, all without the need for retraining the model. We consider a post-disaster environment where an LLM is hosted at the network edge. The UAV can physically approach individual ground sensors. The short, LoS-dominant communication link between a UAV and a ground sensor offers a significant channel gain and supports high-speed data transmission. In this context, employing a UAV for data collection enhances network throughput and extends coverage beyond what terrestrial gateways can achieve[10].
Fig. 3 illustrates a typical UASNETs, where ground sensors are deployed to monitor temperature and humidity in the disaster area. These ground sensors generate environmental data, which is temporarily stored in data queues for later transmission to the UAV. The UAV is deployed to hover over the disaster area and can maneuver close to each ground sensor, leveraging the short-distance LoS communication links to efficiently collect the stored data.
IV-A Path Planning
As depicted in Fig. 3, the UAV hovers over the disaster area where it adjusts its path relative to the requirements of the ground sensors. For example, the UAV must adjust its path based on channel conditions and the queue length of the ground sensors. Moving to the ground sensor with poor channel conditions or a small buffer length for data collection gives rise to packet reception errors or buffer overflow at other ground sensors. Therefore, online path planning of the UAV for data collection is crucial [11].
ML is essential for UAV placement and path planning due to the complex nature of designing UASNETs and the critical requirements in applications such as public safety. Determining the optimal horizontal and vertical placement of the UAV, as well as planning its trajectories to interact with ground or flying objects, requires DRL optimizers that can handle large-scale optimization problems and dynamic environmental factors. In a nutshell, DRL is indispensable for UAV placement and path planning and facilitates complex optimization tasks, supports real-time adaptation to changing conditions, and enhances capabilities for detecting and responding to noncompliant UAV in critical applications like public safety. Path planning involves the optimization of continuous variables. In this work, we simplify the problem by discretizing the action space and restricting the UAV’s motion to a finite number of directions (e.g., north, south, east, and west). This abstraction not only reduces the computational complexity but also makes the problem accessible for LLM-enabled ICL. Starting with a task description depicted in Fig. 2(b), which defines the task and provides in-context demonstrations, the LLM as optimizer generates candidate solutions. These solutions are then implemented in the environment, and their associated costs are computed. Feedback is derived from the latest sensory data, the implemented solution, and the obtained cost. This feedback, along with updated sensory data, is fed back into the LLM. The process repeats iteratively until satisfactory solutions are generated, enabling the LLM to refine its outputs dynamically based on real-world performance.
IV-B Velocity Control
As depicted in Fig. 3, the UAV hovers over the disaster area where it adjusts its velocity based on the requirements of the ground sensors. The slow mobility of a UAV can lead to buffer overflows at the ground sensors, as newly arriving data is not immediately captured by the UAV. The fast mobility of a UAV can harm channel conditions and increase communication failures. Onboard velocity control of the UAV is essential to minimize packet loss due to ground sensor buffer overflows and communication failures[12]. An LLM-enabled ICL framework can be employed to optimize the velocity of the UAV to minimize the overall packet loss. Starting from a task description depicted in Fig. 2(a), which defines the velocity control task and provides in-context examples, the LLM acts as an optimizer to generate candidate velocities. The performance of the candidate velocities is evaluated by calculating a cost metric. The feedback is then shaped by the latest sensor data, velocity, together with the calculated cost. This feedback is fed back to the LLM along with the current sensor data. The LLM iteratively refines its control strategies using this closed-loop process until the system converges to stable, optimal velocity tracking.
IV-C Data Collection Schedule
As depicted in Fig. 3, the UAV hovers over the disaster area, enabling it to approach ground sensors closely and utilize short-range LoS communication links for efficient data collection. However, selecting a specific ground sensor for data collection may result in buffer overflows at other ground sensors if their data queues are near capacity while new data continues to accumulate. Moreover, ground sensors located farther from the UAV typically experience degraded channel conditions, increasing the likelihood of transmission errors. Therefore, an effective data collection schedule is essential to mitigate data queue overflows and reduce communication failures[13]. LLM-enabled ICL is crucial for data collection scheduling in UASNETs, as it facilitates dynamic adaptation to changing environments. By leveraging ICL, the UAV can continuously adjust its data collection scheduling strategies in real- time, selecting the most appropriate sensors. This enables the system to make informed, context-aware decisions without retraining, thereby optimizing data collection efficiency and responsiveness in dynamic or mission-critical scenarios. This ensures efficient, reliable, and autonomous data collection operations, enhancing overall system performance. The relevant task description is depicted in 2(e). This system uses an LLM to optimize data collection scheduling for minimal packet loss by processing real-time sensor inputs (queue lengths and channel conditions) while maintaining constraints that prioritize sensors with high queue lengths (to avoid buffer overflows) and those with good channel conditions (to ensure reliable transmission). The LLM generates an adaptive collection schedule, which is then deployed, recording the resulting system performance (packet loss) and the latest environmental state as feedback. This feedback loop, consisting of the LLM’s decisions, the latest system state, and performance metric, enables continuous improvement by providing information for subsequent planning decisions, allowing the LLM to iteratively refine its policy to minimize packet loss.
IV-D UAV Power Control
We need to optimize energy efficiency and power control in UASNETs because the UAV and ground sensors have limited energy resources, making the optimization of energy and power consumption crucial for the overall performance of the UASNETs. DRL solutions have been proposed to exploit the network’s resources in a near-optimal manner, aiming to maximize energy efficiency effectively. However, these solutions require complex training, and their deployment is not rapid and straightforward. LLM-enabled ICL can be leveraged to optimize the transmit power in UASNETs to maximize overall energy efficiency.
Task description for UAV power control is depicted in 2(c) with a task objective, and the example set represents operational state and environmental context to assist the LLM in inferring optimal power control strategies. This system leverages an LLM to dynamically optimize UAV transmit power for maximum energy efficiency by processing real-time power levels and channel conditions, while adhering to constraints prioritizing sensors with low remaining energy (preventing node failure) and favoring those with strong channel conditions (ensuring reliable transmission). The LLM generates power adjustment commands executed, with the resulting system performance and updated operational state being captured as feedback. This closed-loop mechanism, incorporating the LLM’s power decisions, the latest UAV status, and communication outcomes, enables continuous refinement of power allocation strategies, progressively allowing the UAV to achieve optimal energy efficiency. This approach enables adaptive decision-making without the need for complex training. Such optimization is particularly valuable in public safety scenarios such as emergency response, where efficient power usage directly impacts mission duration and system performance.
(a) Normal and Under attack operation of ICLDC where the jailbreaking attack happens.
(b) Performance with changing the number of ground sensors.
Figure 4: Performance analysis of ICLDC in two scenarios with 10 ground sensors.
IV-E Transmit Power of Ground Sensors
Power control in UASNETs, which involves explicitly managing the transmit power of ground sensors, plays a key role in enhancing system performance by optimizing interference management, signal quality, and resource usage, all while accounting for the limited battery capacities of the ground sensors. This system employs an LLM to optimize ground sensor transmit power for maximal energy efficiency by analyzing current power levels and channel conditions while following constraints that prioritize sensors with high queue lengths (to prevent buffer overflow) and favor good channel conditions. Based on the task description depicted in 2(d), the LLM generates power adjustment recommendations that are implemented, with the resulting system performance and last environmental state being monitored as feedback. This continuous feedback loop, tracking the LLM’s power decisions, the latest sensor data, and transmission outcomes, enables dynamic refinement of power settings, allowing the system to progressively improve energy efficiency while maintaining reliable data transmission and preventing queue overflows. LLM-enabled ICL offers a lightweight alternative to DRL for power control in UASNETs, where it replaces retraining with example-based reasoning and enables real-time decision-making with minimal computational overhead, which is especially well-suited for dynamic, resource-constrained, and safety-critical environments where DRL’s complexity and retraining demands become limiting factors.
Table II: Comparison of Optimization Approaches for Public Safety UAV
Feature DRL Optimizers DM + DRL Hybrid LLM Optimizer (ICL)
Training Requirements Complex model training Improved Model Training No retraining or parameter updates
Deployment Speed Not compatible with urgent requirements Improves real-time performance Rapid adaptability and deployment
Real-World Gap Gaps between simulation and deployment Creates realistic simulation environments Natural language problem expression
Data Dependency High Alleviates data scarcity Problem described in natural language
Adaptability Action-oriented, lacks flexibility Improves adaptability Facilitates rapid task adaptation
Key Advantage State-of-the-art for UASNETs Robust solution for complex scenarios Flexible and generalizable
Key Limitation Slow training and deployment Still requires complex integration Convergence properties unproven
V Case Study: Data Collection Schedule
This section presents a case study focused on data collection scheduling. The objective is to demonstrate how ICL effectively manages data collection scheduling and is vulnerable to jailbreaking attacks. An ICL-based Data Collection Scheduling (ICLDC) framework for UASNETs is proposed in [14]. The proposed ICLDC leverages ICL to optimize the UAV’s emergency data collection schedules. Specifically, the UAV gathers sensory information, such as queue lengths and channel conditions, from ground sensors and transmits this data to an edge-hosted LLM. The LLM is initialized with a structured system prompt that outlines its operational role, constraints, and ethical safeguards, ensuring compliance with safety and security requirements. Upon receiving the environmental data, the LLM generates a detailed task description in natural language, encompassing the mission objective, input parameters, operational rules, illustrative examples, expected outputs, and a feedback mechanism. Based on this context, the LLM infers an optimal data collection schedule, which the UAV subsequently executes. Performance metrics and updated environmental states are continuously monitored and incorporated as feedback into the evolving task description. This feedback is utilized in subsequent decision cycles, enabling the system to adapt dynamically and improve scheduling efficacy over time. Moreover, the paper examines the system’s vulnerability to adversarial manipulation, specifically, jailbreaking attacks, where the task description is maliciously altered to degrade scheduling performance and increase network cost. This highlights the critical need for robust safeguards when deploying LLM-enabled control schemes in mission-critical applications. Fig. 4(a) illustrates the performance of the proposed ICLDC under attack in comparison to the normal scenario. The results clearly show that the attack substantially degrades performance, resulting in higher costs. The attacker employs context-based manipulation to alter the task description within the LLM’s prompt. This deliberate interference misleads the LLM, prompting it to choose suboptimal or potentially harmful actions, which in turn disrupts the overall system behavior. Fig. 4(b) illustrates the performance of ICLDC compared to the under-attack baseline as the number $N$ of ground sensors increases. In general, a larger number of sensors leads to higher costs due to an increased likelihood of buffer overflow. Notably, ICLDC consistently achieves the lowest cost, demonstrating its effectiveness in managing network dynamics.
VI Discussion: From DRL to LLM Optimizers
This section discusses the limitations of DRL algorithms and the need for LLM optimizers like ICL in UASNETs for emergency scenarios. DRL algorithms are considered state-of-the-art optimizers for UASNETs. However, DRL-enabled optimizers tend to be very action-oriented and suffer from challenges such as complex model training and gaps between simulation and real-world deployment, which hinder their rapid deployment in emergencies. In particular, the time-consuming nature of model training and deployment is not compatible with the urgent requirements of public safety situations. What we need are optimizers that offer rapid adaptability and deployment to meet the critical requirements of emergency scenarios.
A promising approach to complement DRL algorithms is the integration of DMs. By combining the data generation capabilities of DMs with the decision-making framework of DRL, this integration improves the adaptability and real-time performance of UASNETs. In addition, studies have shown how DMs can alleviate data scarcity, improve policy networks, and create simulation environments that provide a robust solution for complex UAV communication scenarios[15]. However, even with these advances, there is still a need for lighter-weight solutions that can overcome the challenges of complex model training and the gap between simulation and real-world operations. Rapid adaptability and simplicity of deployment remain key factors for UASNETs in emergencies where time and efficiency are critical.
Recent advancements in prompting techniques have significantly enhanced the performance of LLMs across diverse domains. Leveraging their intrinsic ability to comprehend and generate natural language, LLMs introduce a novel paradigm for addressing optimization problems. Conventional optimization approaches typically require explicit formalization of the problem and the derivation of update rules through specialized solvers. In contrast, LLM-enabled optimization frameworks allow the problem to be expressed in natural language, enabling the model to iteratively propose candidate solutions based on the problem description and previously generated outputs.
The LLM-enabled optimization framework offers several important advantages. First, it facilitates rapid adaptation to different UAV tasks by simply modifying the problem description contained in the prompt. Second, the iterative nature of the process enables continuous refinement of solutions without the need to retrain or change the underlying model parameters. Finally, the optimization process can be effectively customized through prompt engineering by including specific instructions that enforce desired solution properties such as feasibility constraints or performance criteria.
Such ICL capabilities enable LLMs to act as flexible and generalizable optimization agents. This change has the potential to accelerate the development of solutions in UASNETs where explicit mathematical modelling is a challenge. However, further research is needed to accurately quantify the convergence properties and solution quality that can be achieved by LLM-based optimization frameworks.
Overall, while DRL algorithms are state of the art for optimizing UASNETs, they reach their limits in emergency scenarios due to their slow training, complex application and gaps between simulation and reality, making them unsuitable for time-critical situations. DMs improve DRL by generating synthetic data and improving the adaptability of policies, but are still computationally intensive. In contrast, LLM optimizers that use ICL offer a promising alternative. They enable fast, training-free adaptation through natural language prompts, iterative refinement of the solution, and optimization that can be interpreted by humans - important advantages for emergency response. However, LLM-based approaches require further validation of their convergence properties and real-world reliability, suggesting a potential for hybrid systems that combine the robustness of DRL with the flexibility of LLM to meet urgent public safety needs. Table II compares the Optimization Approaches for UASNETs in Emergency Scenarios
VII Most recent trends
This section highlights the future direction of LLM-enabled ICL and the ways to improve the performance and deployment.
VII-1 Enriching Examples using Diffusion Models
Focus on enhancing the generalization capabilities of public safety UAV by addressing the scarcity of high-quality training data. Develop methods to generate diverse and realistic synthetic examples using DMs to augment limited emergency datasets. Further research should explore optimizing DM-based data generation to ensure the quality and relevance of synthetic examples for robust generalization.
VII-2 On-device LLMs for Public Safety UAV
Advance the deployment of LLMs on resource-constrained edge devices, such as UAV platforms, by leveraging recent progress in model compression, efficient architectures, and hardware-software co-optimization. Focus on enabling real-time, localized LLM processing on UAVs to eliminate reliance on remote cloud servers. This will facilitate the creation of personalized, context-aware, and low-latency AI functionalities in UASNETs.
VII-3 Enhanced Robustness Against Adversarial Attacks
Investigate and develop advanced defense mechanisms to mitigate vulnerabilities such as jailbreaking attacks, which manipulate prompts to bypass built-in safety mechanisms and cause LLMs to generate unauthorized or malicious content. Address both white-box and black-box categories of adversarial attacks to ensure the reliability of LLM-enabled ICL in mission-critical scenarios.
VII-4 Extension to Public Safety UAV Swarms
Extend the application of LLM-enabled ICL to public safety UAV swarms, focusing on facilitating coordination among multiple UAVs. Develop methods that enable cooperative task execution in the swarm, leveraging the contextual understanding and adaptive reasoning capabilities of LLMs to improve overall swarm performance and mission success.
VII-5 Real-World Validation and Transferability
Conduct extensive field trials to evaluate the performance of LLM-enabled ICL in real-world public safety applications. Explore the transferability of LLM-enabled ICL to other UAV applications, such as agriculture and infrastructure inspection, by repurposing task descriptions and prompts.
VIII Conclusion
This paper advocates integrating LLM-enabled ICL into public safety UAV to address key operational functions, such as path planning and velocity control. By leveraging ICL, UAVs can dynamically interpret task objectives, adjust behaviors, and make context-aware decisions based on natural language instructions and a few example prompts without the need for retraining. By classifying critical UAV functions and formulating corresponding task descriptions, we demonstrate how LLMs can effectively support path planning, velocity control, data collection scheduling, and power control. The proposed LLM-enabled ICL framework offers a lightweight, adaptable, and scalable approach, setting the stage for new generation of intelligent, edge-enabled public safety UAVs. We envision the application of ICL to public safety UAV swarms, along with its real-world validation and potential use in other domains.